OpenStudy (anonymous):

solve the system of equations below by graphing both equations with a pencil and paper what is the solution? y=-3x+1 y=-x+5

OpenStudy (anonymous):

it would be a lot easier and quicker to solve \[-3x+1=-x+5\] for \(x\) in three steps then find \(y\)

OpenStudy (anonymous):

A. (3,2) B. (0,5) C. (-2,7) C. (-1,4)

OpenStudy (anonymous):

if you don't feel like solving that equation for \(x\) check your answer choices and see which one works solving is probably easier though

OpenStudy (anonymous):

A. (3,2) B. (0,5) C. (-2,7) C. (-1,4)

OpenStudy (usukidoll):

just graph y=-3x+1 (one in blue ink) y=-x+5 (one in red ink) letting x be any number you choose and the point of intersection of these two lines will be the answer.